What is the volume in ml of 20 vol H2O2 required to completely react in 500 ml of 0.02M acidified KMnO4 solution ?

Grade:11

1 Answers

Mayank
4 years ago
 
using the formulae 20vol =5.6*N where N=normality of H2O2
we get N=20/5.6
now equating the equivalents of both the half rxns we get 
20*V/5.6=0.02*5*500 we get V=14ml     // note: valency factor of KMnO4 is 5